The base and height of the triangle are in the ratio 5:3 . If the area of the triangle is 270mm^2; find its base and height.
Answer by lwsshak3(11628)

The base and height of the triangle are in the ratio 5:3 . If the area of the triangle is 270mm^2; find its base and height.
***
let b=base
let h=height
..
b/h=5/3
3b=5h
b=5h/3
area of a triangle=1/2*base*height
1/2*b*h=270
(1/2)(5h/3)(h)=270
5h^2/6=270
5h^2=6*270
h^2=6*270/5=324
h=√324
h=18
b=5h/3=30
base=30 mm
height=18 mm